### v1

# Fixed Points and Universality Classes in Coupled Kardar-Parisi-Zhang Equations

Abstract. Studied are coupled KPZ equations with three control parameters $X,Y,T$. These equations are used in the context of stretched polymers in a random medium, for the spacetime spin-spin correlator of the isotropic quantum Heisenberg chain, and for exciton-polariton condensates. In an earlier article we investigated merely the diagonal $X=Y$, $T=1$. Then the stationary measure is delta-correlated Gaussian and the dynamical exponent equals $z=\tfrac{3}{2}$. We observed that the scaling functions of the dynamic correlator change smoothly when varying $X$. In this contribution, the analysis is extended to the whole $X$-$Y$-$T$ plane. Solutions are stable only if $XY\geq 0$. Based on numerical simulations, the static correlator still has rapid decay. We argue that the parameter space is foliated into distinct universality classes. They are labeled by $X$ and consist of half-planes parallel to the $Y$-$T$ plane containing the point $(X,X,1)$.
1   Introduction
The KPZ equation was introduced by Kardar, Parisi, and Zhang [1] as a model for growing surfaces. The equation governs the motion of a height function, $h(x,t)$, where $x\in\mathbb{R}$ is the substrate coordinate and $t\geq 0$ time. In dimensionless units the KPZ equation is written as
$\partial_{t}h(x,t)=\lambda(\partial_{x}h(x,t))^{2}+\tfrac{1}{2}\partial_{x}^{2}h(x,t)+\xi(x,t),$ (1.1)
where $\xi(x,t)$ is standard spacetime Gaussian white noise. The mean zero time-stationary measure is known to be standard spatial white noise, which immediately implies the wandering exponent $\chi=\tfrac{1}{2}$. From the sum rule $\chi+z=2$ [2, 3], one infers the dynamical exponent $z=\tfrac{3}{2}$. On such scale there are universal distribution functions. For example, for flat initial conditions the height at the origin has fluctuations as
$h(0,t)-\tfrac{1}{3}\lambda^{3}t\simeq\mathrm{sgn}(\lambda)\big{(}|\lambda|t/4\big{)}^{1/3}\xi_{\mathrm{TW}}$ (1.2)
for large $t$. The random variable $\xi_{\mathrm{TW}}$ is distributed according to GOE Tracy-Widom [4, 5]. For other models in the KPZ universality class the coefficient $\lambda$ will be modified, but $\xi_{\mathrm{TW}}$ remains unaltered. Up to scale factors, the KPZ equation has a single universality class.
Our interest is the same equation with several components. Using the Einstein summation convention and maintaining for the moment all model parameters, the generalization reads
$\partial_{t}h_{\alpha}=G^{\alpha}_{\beta\gamma}(\partial_{x}h_{\beta})(\partial_{x}h_{\gamma})+\tfrac{1}{2}D_{\alpha\beta}\partial_{x}^{2}h_{\beta}+B_{\alpha\beta}\xi_{\beta},$ (1.3)
$\alpha,\beta,\gamma=1,...,n$. Here the $n\times n$ matrices $G^{\alpha}$, $D$, and $B$ are given constants and the noise components $\xi_{\alpha}$ are independent. In this note we will consider only $n=2$. Even then there are still many model parameters. A physically natural simplification arises from studies in the context of nonlinear optics, which have the goal to experimentally confirm KPZ physics. One favoured system are two coupled exciton-polariton condensates. Their effective description through coupled KPZ equations is extensively covered in the recent article [6]. Since the condensates are physically indistinguishable, Eq. (1.3) has to be invariant under the interchange $h_{1}\leftrightarrow h_{2}$. Using this symmetry, it can be shown [6] that the $h_{\alpha}$-fields, rotated by $\pi/4$ and appropriately rescaled, are governed by
$\partial_{t}h_{1}$ $=2X(\partial_{x}h_{1})(\partial_{x}h_{2})+\tfrac{1}{2}T\partial_{x}^{2}h_{1}+\sqrt{T}\xi_{1},$ (1.4)
$\partial_{t}h_{2}$ $=Y(\partial_{x}h_{1})^{2}+(\partial_{x}h_{2})^{2}+\tfrac{1}{2}\partial_{x}^{2}h_{2}+\xi_{2},$
where $X,Y$ are real parameters and $T>0$. Details are provided in A. These coupled KPZ equations first appeared in a study of Ertaş and Kardar [7], who investigated the dynamics of stretched polymers immersed in a random medium. More recently the same equations reappeared in the study of sliding particles on a fluctuating landscape [8, 9]. In the context of quantum spin chains of great interest is the equilibrium spacetime correlator of the isotropic Heisenberg chain [10]. In simulations, also in experiments, it was observed that this correlator has the dynamical exponent $z=\tfrac{3}{2}$ [11]. Moreover the shape turned out to be in good agreement with the stationary scaling function of Eq. (1.1). To explain such findings a pair of coupled KPZ equations has been derived in [10], which equals (1.4) upon setting $T=1$.
In (1.1), (1.3), and (1.4) we have omitted a term linear in $\partial_{x}h_{\alpha}$. For the KPZ equation ($n=1$) this term is removed by a Galilei transformation. For $n\geq 2$ the linear term would be of the form $F_{\alpha\beta}\partial_{x}h_{\beta}$. In general, the matrix $F$ is nondegenerate and the linear term can no longer be transformed away. This results in $n$ modes separating ballistically in time. Such case of nondegenerate eigenvalues has been studied in a variety of models [12, 13, 14, 15, 16, 17]. As worked out in great detail [18], the universal features can then be reduced to the single component case. In accordance with the applications mentioned already, we set $F=0$ and require zero initial slope. This implies that all modes have zero velocity and thus interact strongly.
For the KPZ equation the mean zero time-stationary measure can be computed explicitly. Switching to two components, the mean zero time-stationary measure is not known. Only under the condition $Y=X$ the $G^{\alpha}$ matrices are cyclic [18, 19], implying that the steady state consists of two independent copies of spatial Gaussian white noise of strength $1$. To stress, the dynamics could be unstable and no such steady state would exist. However, if a mean zero time-stationary measure does exist and has rapid decay of correlations, then the analogue of the wandering exponent still takes the value $\chi=\tfrac{1}{2}$. The heuristic arguments for the sum rule $\chi+z=2$ rely only on the observation that diffusion and noise govern the microscopic scale, order of the correlation length, while the nonlinear terms govern the widely separated mesoscopic scale. Such reasoning does not depend on the number of components. Thus, as for the KPZ equation, we conclude that for Eq. (1.4) the dynamical exponent $z=\tfrac{3}{2}$, provided there is steady state with short-ranged correlations, see Section 2.1 for the full argument.
Given the scaling exponent, finer details of universality are encoded in scaling functions. Of course, these depend on the particular observable. For $n=1$, there is a single universality class with an example provided in Eq. (1.2), which refers to height fluctuations in case of flat initial conditions. For $n=2$, we focus on the spacetime dynamic correlator of the two components of the slope field in the steady state. Structurally the equations become richer, since one has to take into account linear combinations of the fields. Hence the matrix of scaling functions is now considered to belong to the same universality class, if equal up to dilations and rotations. In principle, the scaling functions depend on all control parameters, $X,Y,T$. Within the set of parameters satisfying $\chi=\tfrac{1}{2}$, $z=\tfrac{3}{2}$, the central issue is to understand how such scalings depend on model parameters. To clarify this issue will be the main task of our contribution.
Figure 1: Setting $T=1$, displayed is the anticipated phase diagram in the $XY$ plane. In the entire stable regime, $XY>0$, the dynamical exponent equals $z=\tfrac{3}{2}$. The diagonal, $\{X=Y\}$, is a line of fixed points with distinct scaling functions. The green half-lines constitute the universality classes. As exemplified in blue, starting with parameters on a particular green line the large time scaling function will be equal to the one of the corresponding fixed point, up to nonuniversal coefficients. The colored lozenges and asterisks indicate parameter points at which direct numerical simulations of Eqs. (1.4) are reported, see Section 3. The black lozenge refers to the limit along the diagonal, see Sections 2.6 and 2.8. The half-line $\{X=1\}$ allows for a Cole-Hopf transformation, see Sections 2.4 and 2.5.
As a guide to the reader, for $T=1$, the anticipated phase diagram is shown in Figure 1. The diagonal $\{X=Y\}$ is a line of fixed points. Along this line, scaling functions are smoothly varying in $X$. For general bare parameters $(X,Y)$, in the long time limit the scaling functions converge to the ones of the diagonal point $(X,X)$. In Figure 1, this feature is indicated by the blue lines with arrow pointing towards the fixed point. We argue that the universality classes consist of the half-lines parallel to the $Y$-axis through parameter point $(X,X)$. Distinct half-lines are separate universality classes. For general $X,Y,T$, the universality classes are argued to be labeled by $X$ and to consist of half-planes parallel to the $Y$-$T$ plane containing the point $(X,X,1)$.
In an earlier contribution [19], we studied universality along the diagonal $\{X=Y\}$ with $T=1$ and confirmed that the scaling functions indeed change smoothly in their dependence on $X$. This work expands the analysis to the full $XY>0$ part of the plane and to general $T>0$. Our text is divided into a theoretical and numerical part. In Section 2 discussed are stability, time-stationary measures, Cole-Hopf transformation, the link to directed polymers, dynamical correlator, cyclicity, and the full phase diagram. Added is a subsection on the boundary points $X=0,Y>0$ and $Y=0,X>0$. Interestingly enough, their behavior foreshadows the transition to an unstable regime defined by $XY<0$. In Section 3 displayed and commented upon are direct numerical simulations at the parameter points marked in Figure 1 and beyond. Finally, in the discussion section we compare our findings with the results based on the one-loop RG analysis in [7], [6]. The heavily used rescalings are detailed in A
2   Two coupled stochastic Burgers equations
For our purposes it will be convenient to switch from height functions to slopes defined as $\phi_{\alpha}=\partial_{x}h_{\alpha}$. Then the two component system under study is
$\partial_{t}\phi_{1}$ $=\partial_{x}\big{(}2X\phi_{1}\phi_{2}+\tfrac{1}{2}T\partial_{x}\phi_{1}+\sqrt{T}\xi_{1}\big{)},$ (2.1)
$\partial_{t}\phi_{2}$ $=\partial_{x}\big{(}Y\phi_{1}^{2}+\phi_{2}^{2}+\tfrac{1}{2}\partial_{x}\phi_{2}+\xi_{2}\big{)},$
where $X,Y$ are arbitrary real parameters and $T>0$. The noise $\xi_{\alpha}(x,t)$ is standard spacetime Gaussian white noise with covariance
$\langle\xi_{\alpha}(x,t)\rangle=0,\quad\langle\xi_{\alpha}(x,t)\xi_{\beta}(x^{\prime},t^{\prime})\rangle=\delta_{\alpha\beta}\delta(x-x^{\prime})\delta(t-t^{\prime}).$ (2.2)
These are two stochastic conservation laws with quadratic nonlinearity, hence of stochastic Burgers type.
2.1   Rescaling and $(1:2:3)$-scaling
In Eq. (1.3) every term has a distinct strength parameter and formulas tend to be unwieldy. Thus it is more convenient to switch to dimensionless units. For this purpose one chooses $\ell$ as unit of space, $\tau$ as unit of time, and the two field amplitudes as $h_{\alpha}(x,t)=a_{\alpha}\tilde{h}_{\alpha}(x/\ell,t/\tau)$. Then the form of the equation is not changed, but $a_{\alpha},\ell,\tau$ can be used to reduce the number of free parameters. For $n=2$ and symmetry $h_{1}\leftrightarrow h_{2}$, the result is stated in Eqs. (1.4), equivalently in Eqs. (2.1) and (2.2), which have 3 dimensionless parameters. The respective computation is explained in A.
A distinct argument concerns the separation between microscopic and mesoscopic scales. For such purpose we introduce the dimensionless parameter $\varepsilon>0$ and investigate the limit $\varepsilon\to 0$. Conventionally used is the scaled height function [3], which is given by
$h_{\varepsilon,\alpha}(x,t)=\varepsilon^{\chi}h_{\alpha}(\varepsilon^{-1}x,\varepsilon^{-z}t).$ (2.3)
Here $x,t$ are of order 1, $\chi$ is the wandering exponent, and $z$ the dynamical exponent. Using the results from A, one arrives at
$\partial_{t}h_{\varepsilon,1}$ $=2\varepsilon^{2-z-\chi}X(\partial_{x}h_{\varepsilon,1})(\partial_{x}h_{\varepsilon,2})+\tfrac{1}{2}\varepsilon^{2-z}T\partial_{x}^{2}h_{\varepsilon,1}+\varepsilon^{{(2\chi-z+1)/2}}\sqrt{T}\xi_{1},$ (2.4)
$\partial_{t}h_{\varepsilon,2}$ $=\varepsilon^{2-z-\chi}Y(\partial_{x}h_{\varepsilon,1})^{2}+\varepsilon^{2-z-\chi}(\partial_{x}h_{\varepsilon,2})^{2}+\tfrac{1}{2}\varepsilon^{2-z}\partial_{x}^{2}h_{\varepsilon,2}+\varepsilon^{{(2\chi-z+1)/2}}\xi_{2}.$
To have a finite limit, neither zero nor infinity, for $h_{\varepsilon,\alpha}$ requires the sum rule
$\chi+z=2.$ (2.5)
If the steady state has rapidly decaying correlations, then $\chi=\tfrac{1}{2}$ and hence
$z=\tfrac{3}{2}.$ (2.6)
The diffusion is order $\varepsilon^{\frac{1}{2}}$ and the noise strength order $\varepsilon^{\frac{1}{4}}$. As a consequence, the mesoscopic scale is characterized by height : space : time having the exponents $\tfrac{1}{2}:\tfrac{2}{2}:\tfrac{3}{2}$, which explains the notion of $1:2:3$ scaling.
2.2   Stability
Diffusion and noise is stable but the nonlinearity could generate instabilities. The simplest textbook example reads
$\partial_{t}\phi_{1}=\partial_{x}\phi_{2},\quad\partial_{t}\phi_{2}=-\partial_{x}\phi_{1},$ (2.7)
which amounts to
$\partial_{t}^{2}\phi_{1}=-\partial_{x}^{2}\phi_{1},\quad\partial_{t}^{2}\phi_{2}=-\partial_{x}^{2}\phi_{2},$ (2.8)
which is the wave equation with the wrong sign.
To investigate this issue in our context we start from the general system
$\partial_{t}\phi_{\alpha}+\partial_{x}J_{\alpha}(\phi_{1},...,\phi_{n})=0,$ (2.9)
which can be written in quasi-linear form as
$\partial_{t}\phi_{\alpha}+A_{\alpha\beta}(\vec{\phi})\partial_{x}\phi_{\beta}=0,$ (2.10)
$\alpha,\beta=1,\dots,n$. The matrix $A$ is the flux Jacobian, which depends on $\vec{\phi}=(\phi_{1},\dots,\phi_{n})$. The system is called hyperbolic, if the eigenvalues of $A$ are real for all values of $\vec{\phi}$. If the system is not hyperbolic, then there are solutions growing exponentially fast. In our case, $n=2$ and $J_{1}=-2X\phi_{1}\phi_{2}$, $J_{2}=-Y\phi_{1}^{2}-\phi_{2}^{2}$ resulting in
$A=-2\begin{pmatrix}X\phi_{2}&X\phi_{1}\\ Y\phi_{1}&\phi_{2}\end{pmatrix}.$ (2.11)
The eigenvalues of $A$ satisfy a quadratic equation with discriminant
$\Delta=(X-1)^{2}\phi_{2}^{2}+4XY\phi^{2}_{1}.$ (2.12)
To have $\Delta\geq 0$ requires
$XY\geq 0.$ (2.13)
Therefore in the quadrants II and IV of the $X$-$Y$ plane, those defined by $XY<0$, the solutions of (2.1) exhibit instabilities. Thus only for $XY>0$ we can expect to have a time-stationary measure. As a consequence only quadrants I and III will be considered.
Hyperbolicity is a strong condition, since it has to hold for arbitrary field configurations. There could be interesting phenomena in the unstable regime. In fact, as explained in [6], for polariton-exciton systems $\phi_{\alpha}$ refers to a phase and hence is bounded by definition. Still the phenomenology in the unstable region is very different from stable KPZ.
2.3   Dynamic correlator
To study scaling functions there is a very wide range of options. In this article we give preference to the time-stationary two-point correlator. In a simulation one runs the system with periodic boundary conditions until it has reached stationarity. This state is now regarded as random initial conditions and one samples the dynamic spacetime correlator
$S_{\alpha\beta}(x,t)=\langle\phi_{\alpha}(x,t)\phi_{\beta}(0,0)\rangle,$ (2.14)
[…]
$C_{\alpha\beta}(x)=\langle\phi_{\alpha}(x)\phi_{\beta}(0)\rangle=S_{\alpha\beta}(x,0)$ (2.15)
and the static susceptibility matrix reads
[…]
Note that the dynamics is stochastically invariant under the transformation $\phi_{1}(x,t)$ to $-\phi_{1}(x,t)$, which yields the simplifications
$S_{12}(x,t)=0=S_{21}(x,t),\quad C_{12}(x)=C_{21}(x)=0,\quad\mathsfit{C}_{12}=\mathsfit{C}_{21}=0.$ (2.17)
Therefore listed are only the diagonal entries of the matrices $S,C,\mathsfit{C}$.
In the simulations reported in Section 3, we observe good convergence to stationarity and a rapid decay to 00 of $C_{11}(x)$ and $C_{22}(x)$. Therefore, by the sum rule $\chi+z=2$, we are assured that the dynamical exponent equals $z=\tfrac{3}{2}$ and the scaling hypothesis reduces to
$S_{\alpha\alpha}(x,t)\simeq t^{-\frac{2}{3}}g_{\alpha}(t^{-\frac{2}{3}}x)$ (2.18)
for sufficiently large $(x,t)$. $g_{\alpha}$ is the scaling function. In principle $g_{\alpha}$ depends on $(X,Y,T)$. Hence more explicitly we should write $g_{\alpha,(X,Y,T)}$. Now let us consider two parameters $(X,Y,T)$ and $(X^{\prime},Y^{\prime},T^{\prime})$. If $g_{\alpha,(X,Y,T)}$ and $g_{\alpha,(X^{\prime},Y^{\prime},T^{\prime})}$ differ only by a scale factor, then $(X,Y,T)$ and $(X^{\prime},Y^{\prime},T^{\prime})$ are in the same universality class. If this is not the case, they are in distinct classes.
In general, as discussed in [19], the correlator $S_{\alpha\beta}$ is a full matrix and so is the respective scaling matrix $g_{\alpha\beta}$. Two parameter values are in the same universality class, if the corresponding scaling matrices are related by rotation and dilation. In our case these matrices are diagonal, which fixes the frame, and thus only dilations have to be considered. So our investigations are focused to the question of how the three-dimensional $X$-$Y$-$T$ space with $XY>0$, $T>0$ is partitioned into universality classes.
2.4   Cole-Hopf transformation
As discovered in [7] and more concisely discussed in [20], only upon setting $T=1$, scanning the half-line $\{(1,Y,1),Y>0\}$ allows for a Cole-Hopf transformation and thereby a link to the stochastic heat equation. One starts from the matrix
$R=2\begin{pmatrix}\sqrt{Y}&1\\ -\sqrt{Y}&1\end{pmatrix}$ (2.19)
and defines the transformed fields as
$\tilde{h}=Rh.$ (2.20)
Then the transformed Eq. (1.4) becomes
$\partial_{t}\tilde{h}_{\alpha}=\tfrac{1}{2}(\partial_{x}\tilde{h}_{\alpha})^{2}+\tfrac{1}{2}\partial_{x}^{2}\tilde{h}_{\alpha}+\tilde{\xi}_{\alpha}.$ (2.21)
This looks like magically the system has been decoupled. But the interaction is hidden in the noise term which is still spacetime Gaussian white noise but correlated as
$\langle\tilde{\xi}_{\alpha}(x,t)\tilde{\xi}_{\beta}(x^{\prime},t^{\prime})\rangle=4\big{(}(1+Y)\delta_{\alpha\beta}+(1-Y)(1-\delta_{\alpha\beta})\big{)}\delta(x-x^{\prime})\delta(t-t^{\prime}).$ (2.22)
For $Y=1$ the system becomes indeed decoupled into two independent KPZ equations, which can be seen also directly from (1.4) by a $\pi/4$ rotation.
The Cole-Hopf transformation is defined by
$Z_{\alpha}=\exp(\tilde{h}_{\alpha}),$ (2.23)
which yields the two-component stochastic heat equation
$\partial_{t}Z_{\alpha}=\tfrac{1}{2}\partial_{x}^{2}Z_{\alpha}+\tilde{\xi}_{\alpha}Z_{\alpha}.$ (2.24)
This is the imaginary time Schrödinger equation with a random potential. The path integral solution can be written as a directed polymer moving forward in time and subject to the spacetime random potential $\tilde{\xi}_{\alpha}(x,t)$. This potential is correlated, but the correlations are in force only when the two polymers cross.
To exploit this connection for the dynamic correlator one faces the difficulty that neither the actual nor the transformed time-stationary measure is known. A more accessible observable would be to start from flat initial conditions, $h_{\alpha}(x,0)=0$. Then the transformed initial conditions are also flat and one arrives at two point-to-line directed polymers. It seems plausible that crossings occur only for a negligible fraction of the total time $t$. Therefore the height fluctuations of $\tilde{h}_{\alpha}$ should converge to two independent GOE Tracy-Widom distributions, denoted by here $\vartheta_{\alpha}$. Inverting $R$, up to constant multiplicative factors, the height fluctuations of $h_{1}$ are given by the sum $(\vartheta_{1}+\vartheta_{2})/\sqrt{2}$ and of $h_{2}$ by the difference $(\vartheta_{1}-\vartheta_{2})/\sqrt{2}$. But these are exactly the height fluctuations at $Y=1$. Thus the Cole-Hopf line $\{(1,Y,1),Y>0\}$ is a single universality class with fixed point $(1,1,1)$. This conclusion is further confirmed by [6], where direct numerical simulations are reported at $(1,2,1)$ for flat initial conditions and independent GOE statistics is observed. In addition, in Section 3 we report on numerical simulations of the dynamic correlator at $(0.4,1,1)$ and $(1,1.2,1)$.
2.5   Mapping to directed polymers
Our crossing discussion is somewhat vague. But it can be made more precise for a very particular discretization of Eq. (2.24), known as semi-discrete directed polymer [21]. In this model spacetime is $\mathbb{R}\times\mathbb{Z}$ with coordinates $t,j$. The polymers are piecewise constant up-right paths which are denoted by $\pi(t)$ taking values in $\mathbb{Z}$, $\pi$ reminding of polymer/path. For uniform weight this corresponds to the asymmetric discretization of $\partial_{x}^{2}$, namely to the rate equation
$\frac{d}{dt}Z_{j}(t)=Z_{j-1}(t)-Z_{j}(t),$ (2.25)
which is a random walk with jumps only to the right. Subtracting drift corresponds to a rotation by $\pi/4$ and thereby yields the spacetime employed in Eq. (2.24). In such frame the random walk is symmetric with diffusive spreading.
To define random weights, one assigns independently to each level line a standard white noise $\xi_{j}(t)$. If the path $\pi(t)$ stays on level $j$ over the time interval $[t_{1},t_{2}]$, the path is assigned the added up the “energy”
[…]
where $\{B_{j}(t)\}$ is a collection of independent standard Brownian motions. The energy $E(\{\pi(t)\})$ of the entire path $\pi(t)$ is simply the sum over all level line energies. Between the initial time $t_{0}$ and the final time $t_{n}$, the path has $n-1$ jump times $t_{0}<t_{1}\dots<t_{n}$. Then
$E(\{\pi(t)\})=\sum_{j=1}^{n}\big{(}B_{j}(t_{j})-B_{j}(t_{j-1})\big{)}$ (2.27)
and the path weight is given by $\exp[E(\{\pi(t)\})]$.
An example of a statistical observable would be the distribution of the energy for paths starting at $(0,0)$ and ending at the point $(N,t)$ in the limit $N\to\infty$ and $t=cN$ with fixed $c>0$. Here energy is the analogue of the height in the KPZ equation. Since this model is in the KPZ universality class, the energy should be linear in $t$ plus fluctuations of order $t^{1/3}$ distributed according to GOE Tracy-Widom. A complete proof is available for stationary initial data for which Tracy-Widom GOE is replaced by the Baik-Rains distribution [22].
The case of interest are two directed polymers. So each object is enlarged by the index $\alpha$, the polymers $\pi_{\alpha}(t)$, jump times $t_{j,\alpha}$, and the noise $B_{j,\alpha}$. As in (2.24) the interaction comes from correlated noise, compare with (2.22). The only modification of the energy occurs when both paths overlap, i.e. $\pi_{1}(t)=j=\pi_{2}(t)$ over some time span $t\in[t_{1},t_{2}]$. There are various options. One possibility is to leave the energy of polymer 2 unchanged. Hence let us assume that in some interval of jump times $[t_{j-1,1},t_{j,1}]$ for polymer 1, it overlaps with polymer 2. Then the weight $B_{j,1}(t_{j,1})-B_{j,1}(t_{j-1,1})$ is modified to $Y(B_{j,1}(t_{j,1})-B_{j,1}(t_{j-1,1}))$. With this imposed correlation, the total energy of interacting polymers is then
$E(\{\pi_{1}(t)\},\{\pi_{2}(t)\})=E(\{\pi_{1}(t)\})+E(\{\pi_{2}(t)\})+(Y-1)\sum_{j=1}^{n}{\!{}^{*}}\big{(}B_{j,1}(t_{j})-B_{j,1}(t_{j-1})\big{)},$ (2.28)
where * indicates that the sum is only over time intervals when path 1 overlaps with path 2.
So for $Y=1$ one has two independent copies. But for $Y\neq 1$ polymer 2 has exactly the same statistics as a single polymer, while polymer 1 has an energy modified through overlaps. For large $N$ the overlap events should become negligible. If so, the energy of each polymer behaves linearly in time with fluctuations of size $t^{1/3}$. Flat initial conditions for $h_{\alpha}(x,0)$ map to two directed polymers both starting at $(0,0)$ and ending somewhere on the line $j+t=N$. In the large $N$ limit, the energies of the two polymers, thus $h_{1}(0,t),h_{2}(0,t)$, have a leading term linear in $t$ and noise components which become independent and are distributed according to GOE Tracy-Widom.
2.6   Cyclicity and fixed points
For a short moment we return to the general case (1.3), setting $n=2$. As noted in [18] and further elaborated by [20, 23, 24] for a particular parameter choice one can still compute the time-stationary measure. The linear part has a Gaussian stationary measure with covariance $C_{\alpha\beta}(x)=\delta(x)\mathsfit{C}_{\alpha\beta}$, where $\mathsfit{C}$ is the solution of
$\tfrac{1}{2}\big{(}D\mathsfit{C}+\mathsfit{C}D^{\mathrm{T}}\big{)}=BB^{\mathrm{T}}$ (2.29)
with $(\cdot)^{\mathrm{T}}$ denoting the transpose of a matrix. We define
$\hat{G}^{\alpha}=(\mathsfit{C}^{-1})_{\alpha\beta}G^{\beta}.$ (2.30)
Then $G^{\alpha}$ are called cyclic with respect to $\mathsfit{C}$, if they are of the form
$\hat{G}^{1}=\begin{pmatrix}a&b\\ b&c\\ \end{pmatrix},\qquad\hat{G}^{2}=\begin{pmatrix}b&c\\ c&d\\ \end{pmatrix}$ (2.31)
with arbitrary coefficients $a,b,c,d$. If $G^{\alpha}$ are cyclic, the full nonlinear equation (1.3) has the same invariant measure as the linear part of the equation.
In our specific case
$G^{1}=\begin{pmatrix}0&X\\ X&0\\ \end{pmatrix},\quad G^{2}=\begin{pmatrix}Y&0\\ 0&1\\ \end{pmatrix},\quad\mathsfit{C}=\begin{pmatrix}1&0\\ 0&1\\ \end{pmatrix}.$ (2.32)
The matrices $G^{\alpha}=\hat{G}^{\alpha}$ are cyclic with respect to $\mathsfit{C}=1$ only if
$Y=X$ (2.33)
Figure 2: Spacetime correlations for $X=Y=-1$ and $X=Y\to\infty$, setting $T=1$. For $(-1,-1,1)$, plotted are the rescaled spacetime correlations (a) $S_{11}$ and (b) $S_{22}$ at different times $t=50,100,200,400$. For $(\infty,\infty,1)$ plotted are rescaled spacetime correlations (a) $S_{11}$ and (b) $S_{22}$ at different times $t=150,300,600$. The simulation is based on Eq. (3.1). We set $\mathsfit{b}=3$ and average over $10^{5}$ independent realizations. The dynamic scaling exponent is $0.67$.
and the coupled Burgers equations read
$\partial_{t}\phi_{1}$ $=\partial_{x}\big{(}2X\phi_{1}\phi_{2}+\tfrac{1}{2}T\partial_{x}\tilde{\phi}_{1}+\sqrt{T}\xi_{1}\big{)},$ (2.34)
$\partial_{t}\phi_{2}$ $=\partial_{x}\big{(}X\phi_{1}^{2}+\phi_{2}^{2}+\tfrac{1}{2}\partial_{x}\phi_{2}+\xi_{2}\big{)}.$
Given cyclicity, the universality classes depend on $(X,X,T)$. These classes might not be disjoint. In [19] we tried to clarify this issue by direct numerical simulations in case $T=1$. In Fig. 2 displayed is the scaled dynamic correlator for the parameters $(-1,-1,1)$ and also for $(\infty,\infty,1)$, which is the limit $X\to\infty$ under the constraint $Y=X$, see Section 2.8. In Fig. 2 plotted are the scaling functions $g_{1,(X,X,1)}$ and $g_{2,(X,X,1)}$, which are referred to as peak 1 and 2. At $X=1$ one observes that both peaks are equal to the exact KPZ scaling function. As $X$ is increased, peak 1 becomes fatter and peak 2 taller under constraint of having area $1$ under the peak. The most pronounced anisotropy occurs at $X=\infty$, at which the ratio (peak top 2/peak top 1) $\simeq 2$. Near the top, peak 1 is considerably flatter than the KPZ peak. For $10<X<$ 1 the order of peak heights is reversed. For $X<0$, we have only a single data point at $X=-1$, at which the peaks have equal height. Qualitatively the order of peaks is reflected at the $Y$ axis, but there is no strict reflection symmetry.
Variations in $X$ result in moderate changes of the scaling functions. Further support comes from the cyclic parameter points $(1,1,1)$ in Figure 3 and $(2,2,1)$ in Figure 4. Taking all available evidence into account, we conclude that $g_{1,(X,X,1)}$ and $g_{2,(X,X,1)}$ are non-trivially modified when scanning $X$. Each one of them thus constitutes a distinct universality class.
For this study, in addition we investigate the dependence on $T$ and simulated the parameter points $(2,2,0.5)$ and $(2,2,2)$. As discussed in Section 3, the value of $T$ results only in non-universal coefficients. Hence we expect that a universality class consists precisely of a half-plane parallel to the $Y$-$T$-plane passing through the point $(X,X,1)$.
2.7   The full phase diagram
The task is to study the coupled noncyclic Burgers equation (2.1),
$\partial_{t}\phi_{1}$ $=\partial_{x}\big{(}2X\phi_{1}\phi_{2}+\tfrac{1}{2}T\partial_{x}\phi_{1}+\sqrt{T}\xi_{1}\big{)},$ (2.35)
$\partial_{t}\phi_{2}$ $=\partial_{x}\big{(}Y\phi_{1}^{2}+\phi_{2}^{2}+\tfrac{1}{2}\partial_{x}\phi_{2}+\xi_{2}\big{)}.$
The steady state is no longer $\delta$-correlated and
$\langle\phi_{\alpha}(x)\phi_{\alpha}(x^{\prime})\rangle=C_{\alpha\alpha}(x-x^{\prime}).$ (2.36)
As our standing assumption, confirmed by simulations, $C_{\alpha\alpha}(x)$ decays rapidly and has a finite correlation length, denoted by $\eta$. In particular, the susceptibility
[…]
As a precursor of the mesoscopic scale, the field $\phi_{\alpha}$ is coarse-grained over a scale much larger than $\eta$. Denoting the coarse-grained field by $\tilde{\phi}_{\alpha}$, this results in
$\langle\tilde{\phi}_{\alpha}(x)\tilde{\phi}_{\alpha}(x^{\prime})\rangle=c_{\alpha}\delta(x-x^{\prime}).$ (2.38)
As a next step we want to write down an effective coupled KPZ equation from which the universality classes can be deduced. Universal scaling properties are captured by the $1:2:3$ scaling, see Section 2.1. Thereby the solution $\phi_{\alpha}(x,t)$ consists of two widely separated parts, a microscopic and mesoscopic part: the latter is a smooth envelope function of order $1$, still random, which is perturbed by in essence white noise of strength $\varepsilon^{\frac{1}{4}}$. With this reasoning, the operations of coarse-graining and forming products should commute,
$(\phi_{\alpha}\phi_{\beta})\,\tilde{}=\tilde{\phi}_{\alpha}\tilde{\phi}_{\beta},$ (2.39)
up to small errors.
For a patch on the microscopic scale, the nonlinearites can be regarded as being constant and only diffusion plus noise has to be handled. To account for a finite correlation length we construct a phenomenological Gaussian model, for which purpose it is convenient to use Fourier space. Then the diffusion constant, $D_{\alpha}(k)$, is taken as $k$-dependent, while the noise strength is a constant $\sigma_{\alpha}$. In Fourier space the Langevin equation is then given by
$\partial_{t}\hat{\phi}_{\alpha}(k,t)=-\tfrac{1}{2}k^{2}D_{\alpha}(k)\sigma_{\alpha}^{2}\hat{\phi}_{\alpha}(k,t)+\mathrm{i}k\sigma_{\alpha}\hat{\xi}_{\alpha}(k,t),$ (2.40)
where $\sigma_{1}=\sqrt{T}$ and $\sigma_{2}=1$ in the case under study. The Fourier transform of the static correlator is $\hat{C}_{\alpha\alpha}(k)=1/D_{\alpha}(k)$. In real space, $C_{\alpha\alpha}(x)$ decays rapidly and hence $D(k)$ diverges at large $k$ thereby suppressing the corresponding Fourier modes. The susceptibility is $c_{\alpha}=\hat{C}_{\alpha\alpha}(0)=1/D_{\alpha}(0)$. Coarse-graining over scales much larger than $\eta$ amounts to discarding the large $k$ behavior. Switching back to physical space and denoting the coarse-grained fields by $\tilde{\phi}_{\alpha}$ yields the effective Langevin equation
$\partial_{t}\tilde{\phi}_{\alpha}(x,t)=\tfrac{1}{2}c_{\alpha}^{-1}\sigma_{\alpha}^{2}\partial_{x}^{2}\tilde{\phi}_{\alpha}(x,t)+\sigma_{\alpha}\partial_{x}\xi_{\alpha}(x,t).$ (2.41)
In our approximation spatial coarse-graining merely modifies the diffusion constant.
Adding, after coarse graining, the nonlinear terms present in (2.35) back into (2.41) leads to
$\partial_{t}\tilde{\phi}_{1}$ $=\partial_{x}\big{(}2X\tilde{\phi}_{1}\tilde{\phi}_{2}+\tfrac{1}{2}c_{1}^{-1}T\partial_{x}\tilde{\phi}_{1}+\sqrt{T}\xi_{1}\big{)},$ (2.42)
$\partial_{t}\tilde{\phi}_{2}$ $=\partial_{x}\big{(}Y\tilde{\phi}_{1}^{2}+\tilde{\phi}_{2}^{2}+\tfrac{1}{2}c_{2}^{-1}\partial_{x}\tilde{\phi}_{2}+\xi_{2}\big{)}.$
We now return to the true steady state obtained from (2.35) and recall that under coarse-graining the correlator becomes $c_{\alpha}\delta(x-x^{\prime})$, see (2.36) and (2.38). However a much stronger property holds generically for systems away from criticality. Under coarse-graining, the steady state field $\tilde{\phi}_{\alpha}(x)$ is actually Gaussian white noise of strength $c_{\alpha}$. The two components are independent. We have obtained two properties based on disjoint arguments, the effective evolution equation (2.42) and the white noise statistics of the steady state. Following the discussion in Section 2.6 both properties together necessarily imply that the model is cyclic relative to $\mathsfit{C}_{\alpha\beta}=c_{\alpha}\delta_{\alpha\beta}$. Hence
$\frac{1}{c_{1}}X=\frac{1}{c_{2}}Y.$ (2.43)
Since the model is cyclic, under the $1:2:3$ scaling the properties discussed in Section 2.6 are in force.
Our result becomes more transparent through carrying out standard rescalings. First the fields are normalized as
$\check{\phi}_{\alpha}=\frac{1}{\sqrt{c_{\alpha}}}\tilde{\phi}_{\alpha}$ (2.44)
and the relation (2.43) is implemented. Then the dynamics of the new fields is governed by
$\partial_{t}\check{\phi}_{1}$ $=\sqrt{c_{2}}\partial_{x}\big{(}2X\check{\phi}_{1}\check{\phi}_{2}\big{)}+\partial_{x}\big{(}\tfrac{1}{2}c_{1}^{-1}T\partial_{x}\check{\phi}_{1}+c_{1}^{-1/2}\sqrt{T}\xi_{1}\big{)},$ (2.45)
$\partial_{t}\check{\phi}_{2}$ $=\sqrt{c_{2}}\partial_{x}\big{(}X\check{\phi}_{1}^{2}+\check{\phi}_{2}^{2}\big{)}+\partial_{x}\big{(}\tfrac{1}{2}c_{2}^{-1}\partial_{x}\check{\phi}_{2}+c_{2}^{-1/2}\xi_{2}\big{)}.$
A further rescaling, as in Appendix A, with $a_{1}=a_{2}=c_{2}^{-3/2},\ell=c_{2}^{-3},~{}\tau=c_{2}^{-5}$ leads to the equivalent equation
$\partial_{t}\check{\phi}_{1}$ $=\partial_{x}\big{(}2X\check{\phi}_{1}\check{\phi}_{2}\big{)}+\partial_{x}\big{(}\tfrac{1}{2}T_{\rm eff}\partial_{x}\check{\phi}_{1}+\sqrt{T_{\rm eff}}\xi_{1}\big{)},$ (2.46)
$\partial_{t}\check{\phi}_{2}$ $=\partial_{x}\big{(}X\check{\phi}_{1}^{2}+\check{\phi}_{2}^{2}\big{)}+\partial_{x}\big{(}\tfrac{1}{2}\partial_{x}\check{\phi}_{2}+\xi_{2}\big{)},$
where $T_{\rm eff}=(c_{2}/c_{1})T$. It then follows that the bare parameters $(X,Y,T)$ are in the same universality class as the point $(X,X,T_{\rm eff})$.
Physically one would expect that the large scale behavior depends only on the susceptibility, but not separately on diffusion and noise strength. Thus, the points $(X,X,T)$ with arbitrary $T$ should be in same universality class, which is confirmed by the simulations to be discussed in Section 3.3. By convention, we can regard $(X,X,1)$ as the fixed point defining the class. Then our final conclusion is that $(X,Y,T)$ is in the universality class characterized by the fixed point $(X,X,1)$. This feature is partially visualized in Figure 1 by arrows pointing from the bare $(X,Y)$ towards the fixed point $(X,X)$.
Somewhat unexpectedly, our argument yields the novel relation (2.43) which involves only properties of the steady state. Thereby we have acquired a simple test whether our reasoning is valid. Significant deviations would mean that our arguments require further improvements.
2.8   Limiting parameters
Throughout this sub-section we set $T=1$ and discuss the limiting parameters $(0,0,1)$, $(0,Y,1)$, and $(X,0,1)$. The fourth case is the limit $X\to\infty$ with $Y=\kappa X$, $\kappa>0$.
(0,0): The two components are uncoupled and satisfy a linear Langevin equation, which is referred to as Gaussian universality.
(0,Y): The equations of motion are
$\partial_{t}\phi_{1}$ $=\partial_{x}\big{(}\tfrac{1}{2}\partial_{x}\phi_{1}+\xi_{1}\big{)},$ (2.47)
$\partial_{t}\phi_{2}$ $=\partial_{x}\big{(}Y\phi_{1}^{2}+\phi_{2}^{2}+\tfrac{1}{2}\partial_{x}\phi_{2}+\xi_{2}\big{)}.$
Clearly component 1 is Gaussian. For component 2 the noise $\xi_{2}$ is modified to $Y\phi_{1}^{2}+\xi_{2}$, where the two terms are independent. The truncated spacetime covariance of $\phi_{1}^{2}$ reads
$\langle\phi_{1}(x,t)^{2}\phi_{1}(0,0)^{2}\rangle^{\mathrm{c}}=2\langle\phi_{1}(x,t)\phi_{1}(0,0)\rangle^{2}=\frac{1}{\pi t}\exp(-x^{2}/t).$ (2.48)
This term is decaying to zero for long times. Thus the asymptotics of the $\phi_{2}$ field is expected to satisfy KPZ scaling.
(X,0): The equations of motion are
$\partial_{t}\phi_{1}$ $=\partial_{x}\big{(}2X\phi_{1}\phi_{2}+\tfrac{1}{2}\partial_{x}\phi_{1}+\xi_{1}\big{)},$ (2.49)
$\partial_{t}\phi_{2}$ $=\partial_{x}\big{(}\phi_{2}^{2}+\tfrac{1}{2}\partial_{x}\phi_{2}+\xi_{2}\big{)},$
The $\phi_{2}$ field is decoupled and satisfies KPZ scaling. The feedback to the field $\phi_{1}$ is more tricky. It is still a linear Langevin equation, but there is a linear drift with an independent spacetime stationary random strength. The average $\langle\phi_{1}(x,t)\rangle=0$. This indicates that the $\phi_{2}$ field converges to a Gaussian, as in case $(0,0)$.
$(X,\kappa X),\,\,large\,\,\kappa$: Setting $Y=\kappa X$, the common factor $X$ of the nonlinearity can be absorbed by rescaling time. The coefficient in front of $\phi_{2}^{2}$ equals $1/X$. Hence in the limit $X\to\infty$, the equations read
$\partial_{t}\phi_{1}$ $=\partial_{x}\big{(}2\phi_{1}\phi_{2}+\tfrac{1}{2}\partial_{x}\phi_{1}+\xi_{1}\big{)},$ (2.50)
$\partial_{t}\phi_{2}$ $=\partial_{x}\big{(}\kappa\phi_{1}^{2}+\tfrac{1}{2}\partial_{x}\phi_{2}+\xi_{2}\big{)},$
which should be viewed as a coupled stochastic Burgers equation in its own right. For $\kappa=1$, the model is cyclic and hence the time-stationary measure consists of two independent standard spatial white noises. In fact, $\kappa=1$ has been studied already in [19] corresponding to the case with parameter $\lambda=0$. It is for this choice of parameters that one observes the yet strongest deviations from the KPZ scaling function, compare with Fig. 2. Away from this fixed point the time-stationary measure is not known. However (2.43) turns into the prediction
$\kappa=\frac{c_{2}}{c_{1}}.$ (2.51)
For all values of $\kappa$ one expects the scaling functions to converge asymptotically to the ones of the fixed point $\kappa=1$. From the viewpoint of our theory, Eq. (2.50) has a minimal number of parameters and thus would be an interesting testing ground.
3   Direct numerical simulations
3.1   Static and dynamic correlators
Simulated are the coupled Burgers equations
$\partial_{t}\phi_{1}$ $=\partial_{x}\big{(}2\mathsfit{b}X\phi_{1}\phi_{2}+T\partial_{x}\phi_{1}+\sqrt{2T}\xi_{1}\big{)},$ (3.1)
$\partial_{t}\phi_{2}$ $=\partial_{x}\big{(}\mathsfit{b}X\phi_{1}^{2}+\mathsfit{b}\,\phi_{2}^{2}+\partial_{x}\phi_{2}+\sqrt{2}\xi_{2}\big{)}.$
[…]
The finite grid consists of the integers $1\leq j\leq L$ with periodic boundary conditions, where $L=2048$ for our simulations. To distinguish from the continuum, the fields are denoted by $\phi_{\alpha,j}(t)$. In the initial measure the fields $\phi_{\alpha,j}(0)$ are i.i.d. Gaussians with mean 00 and variance $1$. When $X=Y$, this is already the steady state and one samples the dynamic correlator as
$S_{\alpha\alpha}(j-i,t)=\langle\phi_{\alpha,j}(t)\phi_{\alpha,i}(0)\rangle.$ (3.2)
In our simulations the number of independent samples is of the order $10^{4}$.
However when $X\neq Y$, the situation is slightly more involved. First one has to equilibrate the system, through running the dynamics up to some sufficiently long time, here denoted by $t_{\mathrm{eq}}$. We tested equilibration times from $t_{\mathrm{eq}}=500$ to $t_{\mathrm{eq}}=5000$. As a check of stationarity, considered is the equal time correlator
$C_{\alpha\alpha}(j-i,t)=\langle\phi_{\alpha,j}(t)\phi_{\alpha,i}(t)\rangle.$ (3.3)
For $t>t_{\mathrm{eq}}$ this average should remain unchanged. Applying standard rules blindly, the static susceptibility is the defined by
$\sum_{j=1}^{L}C_{\alpha\alpha}(j,t)=\chi_{\alpha}(t).$ (3.4)
Since Eq. (3.1) is of conservation type, $\chi_{\alpha}(t)$ does not depend on time and hence $\chi_{\alpha}(t)=1$. In actual fact, $C_{\alpha\alpha}(j,t)$ has a peak centered at the origin with width of order $1$ and a constant background of amplitude $1/L$. This can be seen more clearly by considering the Fourier transform
$\hat{C}_{\alpha\alpha}(k,t_{\mathrm{eq}})=\sum_{j=1}^{L}\mathrm{e}^{\mathrm{i}kj}C_{\alpha\alpha}(j,t_{\mathrm{eq}}),$ (3.5)
where $k=(2\pi/L)m$, $m=0,...,L-1$. Then $\hat{C}_{\alpha\alpha}(0,t_{\mathrm{eq}})=1$. But the physical susceptibility equals
$c_{\alpha}=\lim_{k\to 0}\hat{C}_{\alpha\alpha}(k,t_{\mathrm{eq}}).$ (3.6)
Of course, $k$ runs only over a grid with spacing $2\pi/L$. So the limit (3.6) is understood in sense of a quadratic fit close to $k=0$. More precisely, for $k\neq 0$, but close to 00, one uses the fit function $f_{\mathrm{fit}}(k)=c_{\alpha}-ak^{2}$, $a>0$.
The proper definition of the dynamic correlator is then
$S_{\alpha\alpha}(j-i,t)=\langle\phi_{\alpha,j}(t_{\mathrm{eq}}+t)\phi_{\alpha,i}(t_{\mathrm{eq}})\rangle+\tfrac{1}{L}(-1+c_{\alpha}).$ (3.7)
Summing $S_{\alpha\alpha}(j-i,t)$ over $j$, the first summand results in $1$ and both summands add up to $c_{\alpha}$, as it should be. Finally, one notes that upon setting
$\hat{\phi}_{\alpha}(k,t)=\frac{1}{\sqrt{L}}\sum_{j=1}^{L}\mathrm{e}^{\mathrm{i}kj}\phi_{\alpha,j}(t),$ (3.8)
the Fourier transform of the equal time correlator reads
$\langle|\hat{\phi}_{\alpha}(k,t)|^{2}\rangle=\hat{C}_{\alpha\alpha}(k,t).$ (3.9)
As a routine control we confirm that the off-diagonal matrix elements of $C_{\alpha\beta}(j,t)$ and $S_{\alpha\beta}(j,t)$ indeed vanish within statistical error bars.
In the introduction, see Figure 1, displayed are the parameter points for which simulations have been carried out. Out of those, the Table below lists two blocks, $X=1$ and $X=2$, each consisting of three parameter points. One is the fixed point, the two others are above and below the fixed point, all at the same value of $X$. The parameter $\mathsfit{b}$ refers to (3.1) and $p_{\alpha}$ is the maximum height of peak $\alpha$. According to theory, the ratios $c_{2}p_{1}/c_{1}p_{2}$ and $c_{2}X/c_{1}Y$ should be independent of $Y$ for given X.
According to the scheme outlined above, simulated are $C_{\alpha\alpha}(j,t_{\mathrm{eq}})$ and $S_{\alpha\alpha}(j,t)$ for the six parameter points from the Table. The dynamical correlator is scaled with the exponent $\tfrac{3}{2}$ and thus provides information on the scaling functions $g_{\alpha}$. For the cyclic parameters along the diagonal, the susceptibilities are $c_{1}=1$, $c_{2}=1$. As an important control check, we confirm that $C_{\alpha\alpha}(j,t)$ is concentrated at the single site $j=0$ of the numerical grid. For non-cyclic parameters, $C_{\alpha\alpha}(j,t_{\mathrm{eq}})$ is broadened and its susceptibility $c_{\alpha}$ is determined by Eq. (3.6). The results are displayed in Figs. 3 and 4.
Fig. $\mathsfit{b}$ $X$ $Y$ Symbol $c_{1}$ $c_{2}$ $c_{2}p_{1}/c_{1}p_{2}$ $c_{2}X/c_{1}Y$
$3$ $4$ $1$ $1$ red lozenge $1.0$ $1.0$ $1.0$ 1.00
$3$ $5$ $1$ $0.4$ magenta asterisk $1.69$ $0.72$ $0.97$ $1.07$
$3$ $4$ $1$ $1.2$ pink asterisk $0.91$ $1.10$ $1.00$ 1.01
$4$ $3$ $2$ $2$ blue lozenge $1.0$ $1.0$ $0.75$ 1.00
$4$ $3$ $2$ $1.1$ blue asterisk $1.30$ $0.82$ $0.73$ 1.15
$4$ $2.5$ $2$ $2.9$ purple asterisk $0.88$ $1.20$ $0.75$ $0.94$
Table 1: Ratios $c_{2}p_{1}/c_{1}p_{2}$ and $c_{2}X/c_{1}Y$ at parameters used in numerical simulations of (3.1). Here $c_{\alpha}$ is the static susceptibility and $p_{\alpha}$ the height at the origin of the peak $\alpha$. There are two universality classes, $X=1$, $X=2$, and the symbols correspond to the ones displayed in Fig. 1. The values of $c_{\alpha}$ are computed using a quadratic fit for $\hat{C}_{\alpha\alpha}(k,t_{\mathrm{eq}})$, compare with (3.6) and explanations below. The ratios $c_{2}p_{1}/c_{1}p_{2}$ and $c_{2}X/c_{1}Y$ are observed to be in reasonable good agreement in either class.
Figure 3: Simulations for parameters $X=1$, $Y=1,0.4,1.2$. Displayed are the steady state static correlators $C_{11}(j)$ and $C_{22}(j)$, (c), (f), (i). On the scale $t^{0.67}$ plotted are the spacetime correlations $S_{11}$, (a), (d), (g), and $S_{22}$, (b), (e), (h), at different times $t=50,100,200$. The parameter $\mathsfit{b}$ is provided in Table 1 and sampled are $10^{4}$ independent realizations.
One first notes that the static correlator, while broadened, decays very quickly to zero, thereby confirming a key assumption of our theory. Secondly within error bars the scaling exponent of $0.67$ is well confirmed throughout. At $(1,1)$ the peaks have the same height and become asymmetric when moving to $(1,0.4)$ and $(1,1.2)$. This is a nonuniversal feature which is compensated by $c_{1},c_{2}\neq 1$. At $(2,2)$ the fixed point has already an intrinsic asymmetry, which is reduced by moving to $(2,1.1)$ and enforced at $(1,2.9)$, again nonuniversal features.
Figure 4: Simulations for parameters $X=2$, $Y=2,1.1,2.9$. Displayed are the steady state static correlators $C_{11}(j)$ and $C_{22}(j)$, (c), (f), (i). On the scale $t^{0.67}$ plotted are the spacetime correlations $S_{11}$, (a), (d), (g), and $S_{22}$, (b), (e), (h), at different times $t=50,100,200$. The parameter $\mathsfit{b}$ is provided in Table 1 and sampled are $10^{4}$ independent realizations.
From our analysis we concluded that the ratios $c_{2}p_{1}/c_{1}p_{2}$ and $c_{2}X/c_{1}Y$ should be independent of $Y$ for given $X$. In the Table listed are the measured values of $c_{\alpha}$ and from Figures 3 and 4 one reads off the value of $p_{\alpha}$. The resulting ratios are in reasonable good agreement with theory. However, based on prior experience, a much stronger criterion is to compare the full scaling functions, the results of which will be reported next.
3.2   Comparison of scaling functions at ${X=1,2}$
For the KPZ equation there is a well understood scaling theory, which predicts the nonuniversal model-dependent coefficients. In numerical simulations, in particular away from integrability [25], it is common experience that the scaling function has already the predicted shape, while the nonuniversal coefficients show still considerable deviations. Therefore when fitting theory with numerics one introduces a time scale as single free parameter. Ideally, for very long times this factor should be $1$. We adopt the same procedure for our two-component system and introduce the free parameter $s$ through the scaled time $st$ and optimize so to have maximal agreement. Since there is only one time, the scale parameter has to be the same for both components. Separately for $X=1$ and $X=2$, the dynamical correlators from Figures 3, 4 are normalized by $1/c_{\alpha}$ and the parameter $s$ is introduced. In Figures 5, 6 the optimal fit is shown and the parameter $s$ is recorded.
Figure 5: Universal scaling function for $X=1$. For the longest available time $t=200$, the six curves from Figure 3 are multiplied by $1/c_{\alpha}$ and the free time scale $s$ is optimized.
Figure 6: Universal scaling function for $X=2$. For the longest available time $t=200$, the six curves from Figure 4 are multiplied by $1/c_{\alpha}$ and the free time scale $s$ is optimized.
Within numerical errors the fit is viewed as convincing. In [19], we reported on a similar comparison between the scaling functions of a two lane-lattice gas and the corresponding two-component KPZ equation. Both models are cyclic and no equilibration step is required. On the other hand, beyond the factor $1/c_{\alpha}$, a rotation by $\pi/4$ had to be implemented. In this case, the coincidence of the two scaling functions is more striking than the one of Figures 5, 6. Presumably, further averaging would improve the agreement.
3.3   Beyond the subspace $\{T=1\}$
To explore $T\neq 1$, we follow the same protocol used when varying $Y$. Direct simulations are carried out for the parameters $(2,2,T)$ with $T=0.5,1,2$. Since the parameters are cyclic the equilibration time $t_{\mathrm{eq}}=0$. To account for the possible non-universal factors when comparing with $T=1$, the scale $s$ is introduced as a free parameter. For the two values $T=0.5,2$ a good fit with the scaling function at $(2,2,1)$ is achieved. To be noted, the scale factor is $s=0.65$ for $T=0.5$ and $s=0.6$ for $T=2$. More simulations are in demand. But at least for the chosen parameter values, a physically natural scenario is confirmed: For the long time behavior locally the strength of the diffusion term and of the noise term enter only through the thereby determined effective susceptibility.
Figure 7: Simulations for parameters $X=2$, $Y=2$ with $T=2,0.5$. On the scale $t^{0.67}$ plotted are the spacetime correlations $S_{11}$, (a), (c), and $S_{22}$, (b), (d), at different times $t=50,100,200$. The parameter $\mathsfit{b}=2$ and sampled are $5\times 10^{4}$ independent realizations. The data are compared with the fixed point parameters $(2,2,1)$ as displayed in top of Fig. 4.
4   Discussions
We return to the coupled equations
$\partial_{t}h_{1}$ $=2X(\partial_{x}h_{1})(\partial_{x}h_{2})+\tfrac{1}{2}T\partial_{x}^{2}h_{1}+\sqrt{T}\xi_{1},$ (4.1)
$\partial_{t}h_{2}$ $=Y(\partial_{x}h_{1})^{2}+(\partial_{x}h_{2})^{2}+\tfrac{1}{2}\partial_{x}^{2}h_{2}+\xi_{2}.$
To analyse their phase diagram, D. Ertaş and M. Kardar [7] worked out already the RG flow in one-loop approximation and obtained flow equations for the 7 parameters as displayed in Eq. (A). In the recent posting [6], H. Weinberger et al. reworked the flow equation, observing that in fact only four coefficients are needed, namely $X,Y,T$ and an additional global factor, denoted by $Z$, in front of the nonlinearity. As shown in A, the number of dimensionless parameters can be even further reduced to three, namely $(X,Y,T)$. Measuring the RG coarse-graining length $\ell$ in units of the length $1/(4\pi\ell)=\sigma_{2}^{2}\Gamma_{2}^{2}/(4\pi D_{2}^{3})$, see A, and in agreement with [7], [6], the following flow equations hold:
$\frac{d}{d\ell}T=\frac{-2X((T-3)X-3TY+Y)}{(1+T)^{2}}-T-XY,$
$\frac{d}{d\ell}X=-\frac{4(X-1)X(X-Y)}{(1+T)^{2}},$
$\frac{d}{d\ell}Y=\frac{Y(X-Y)(2(5T+1)X+(T+1)^{2}Y-4T(T-1))}{T(T+1)^{2}}.$ (4.2)
A priori, the flow equations are valid for arbitrary $(X,Y,T)\in\mathbb{R}^{3}$. Since $T$ is a ratio of diffusion constants, we require $T>0$. In the present context stability is required, which translates to $XY>0$.
Universality classes are labeled by the fixed points of the flow equations. Setting the left hand side of (4) to zero, one arrives at a line of fixed points defined by
$P_{\mathrm{f}}=(X_{\mathrm{f}},X_{\mathrm{f}},T(X_{\mathrm{f}})),\quad T(X)=\tfrac{1}{2}\big{(}\sqrt{(1+X^{2})^{2}+12X^{2}}-1-X^{2}\big{)}$ (4.3)
with $X_{\mathrm{f}}\in\mathbb{R}\setminus\{0\}$. $P_{\mathrm{f}}$ has two stable directions and one neutral direction with eigenvector pointing along the line of fixed points. Starting with general initial data satisfying $XY>0$, $T>0$, in the limit $\ell\to\infty$, a unique fixed point is reached. The universality class labeled by $X_{\mathrm{f}}$ is the basin of attraction of $P_{\mathrm{f}}$ and consists of all flow lines ending up at $P_{\mathrm{f}}$. Thereby the three-dimensional phase space of Eq. (4) is foliated into universality classes, which are two-dimensional surfaces in three-space.
Of interest are also the two border lines $XY=0$, compare with Section 2.8. For initial $(X,0,T)$ the flow converges to the fixed point $(1,0,1)$. This is in agreement with Section 2.8, where it is argued that $h_{1}$ will be in the KPZ universality class, while $h_{2}$ should be Gaussian. However this fixed point has an unstable direction pointing towards $Y<0$. In this sense, the positive $X$-axis is a borderline. In the unstable regime, $XY<0$, there is a single further fixed point given by $(X,Y,T)=(1,-1,1)$, the properties of which are reported in [6]. The initial data $(0,Y,T)$ flow towards the fixed point $(0,0,0)$, which indicates diffusive scaling for both components. The analysis in Section 2.8 suggests that $h_{1}$ is Gaussian and $h_{2}$ has KPZ scaling.
The RG flow equations (4) refer to successively integrating out large momentum scales in one-loop order. Our method is conceptually more primitive, since we only spatially coarse-grain and subsequently rescale. Both approaches agree on the dynamical exponent $z=\tfrac{3}{2}$. As to universality classes, both methods detect the Cole-Hopf line, i.e. the half-line universality class $\{(1,Y,1),Y>0\}$. This seems to be the only choice of control parameters for which there is a proof-like argument, compare with Section 2.6.
Much more surprisingly, the line of fixed points is detected by both methods. In particular, the universality classes are labeled merely by $X_{\mathrm{f}}$. But the surfaces defining the $X_{\mathrm{f}}$ universality classes differ from each other. In our approach the surface is the $Y$-$T$ plane shifted to $X_{\mathrm{f}}$. For the RG approach one has to compute numerically the basin of attraction of the fixed point with label $X_{\mathrm{f}}$. While this can be done, to have an impression we only plot particular sections. We consider fixed $X=1,2,5$, $T=0.5,1,2$ and solve (2.6) with initial data $P_{0}=(X,Y,T)$ allowing for general $Y$. Then $X_{\mathrm{f}}$ is plotted as a function of $Y$, see Figure 8. According to our claim one should have simply the constant function $X_{\mathrm{f}}=X$ independent of $Y,T$. The flow equations (4) make a distinct prediction.
Figure 8: Figure illustrates the RG flowlines (solid lines) for $X=1,2,5$, represented by green, red and blue respectively. The fixed point $X_{\mathrm{f}}$ is plotted versus $Y$ for a range of $T$ values, i.e., $T=0.4,0.6,0.8,1.0,1.2,1.4$. We notice a relatively small spread while varying $T$. The horizontal lines (dashed) are independent of $T$ and display the predictions based on our work.
To compare with dynamical critical phenomena [26], we focus on model B which is characterized by a single conservation law. Away from static criticality the order parameter field has a finite correlation length and time spreading is diffusive. At or very close to criticality, static correlations become long-ranged. As a consequence, the spreading becomes non-diffusive with a dynamical exponent $z\neq 2$. The KPZ equation has a distinct scenario, however. In the parameter regime studied, static correlations decay exponentially, while the dynamical exponent equals $z=\tfrac{3}{2}$ throughout. In this sense the model is critical. The phase diagram refers to the behavior of scaling functions. For two components, we argue that the coupling constant $X$ is relevant, meaning that the scaling functions vary non-trivially with $X$, while the couplings $Y,T$ are irrelevant.
Of course, one would like to have a more quantitave comparison. Discussing such issue, one first should remark that one-loop RG is an approximation, while our scheme is non-perturbative. A second thought is to accumulate further numerical evidence. For the numerically simulated case $X=2$, $Y=1.1,2.9$, and $T=1$, the two predictions are still within numerical error bars. One could try other values for the control parameters. But this involves additional direct simulations with a modest chance of resolving the issue. A much more promising direction is the well-developed functional renormalization of KPZ [27]. Originally its focus was on KPZ in higher dimensions. But the method has been adjusted to one dimension. Predicted is the Fourier transform of the dynamic correlator. The agreement with the exact solution is impressive. Even the small negative dip in the scaling function is reproduced, see Section VI of [27]. Thus an interesting goal for the future is to extend functional RG to coupled KPZ equations in one dimension. As an additional bonus, our approach and also numerical data could be tested against the sophisticated theory [27].
Appendix A Dimensionless units
Let us start from Eq. (1.3) with $n=2$,
$\partial_{t}h_{\alpha}=G^{\alpha}_{\beta\gamma}(\partial_{x}h_{\beta})(\partial_{x}h_{\gamma})+\tfrac{1}{2}D_{\alpha\beta}\partial_{x}^{2}h_{\beta}+B_{\alpha\beta}\xi_{\beta},$ (A.1)
$\alpha,\beta,\gamma=1,2$. Imposing the constraint of invariance under interchanging the two components, the coupling matrices acquire the special form
$G^{1}=\begin{pmatrix}\Gamma_{11}&\Gamma_{12}\\ \Gamma_{12}&\Gamma_{22}\\ \end{pmatrix},\quad G^{2}=\begin{pmatrix}\Gamma_{22}&\Gamma_{12}\\ \Gamma_{12}&\Gamma_{11}\\ \end{pmatrix},\quad D=\begin{pmatrix}d_{1}&d_{2}\\ d_{2}&d_{1}\\ \end{pmatrix},\quad B=\begin{pmatrix}b_{1}&b_{2}\\ b_{2}&b_{1}\\ \end{pmatrix}.$ (A.2)
It is convenient to rotate the equations of motion by $\pi/4$, to say
$\tilde{h}=Rh,\quad R=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-1\\ 1&1\end{pmatrix},$ (A.3)
which yields
$\partial_{t}\tilde{h}_{1}=2\tfrac{1}{\sqrt{2}}(\Gamma_{11}-\Gamma_{22})(\partial_{x}\tilde{h}_{1})(\partial_{x}\tilde{h}_{2})+\tfrac{1}{2}(d_{1}-d_{2})\partial_{x}^{2}\tilde{h}_{1}+(b_{1}-b_{2})\xi_{1},$
$\partial_{t}\tilde{h}_{2}=\tfrac{1}{\sqrt{2}}(\Gamma_{11}+\Gamma_{22}-2\Gamma_{12})(\partial_{x}\tilde{h}_{1})^{2}+\tfrac{1}{\sqrt{2}}(\Gamma_{11}+\Gamma_{22}+2\Gamma_{12})(\partial_{x}\tilde{h}_{2})^{2}$
$\hskip 40.0pt+\tfrac{1}{2}(d_{1}+d_{2})\partial_{x}^{2}\tilde{h}_{2}+(b_{1}+b_{2})\xi_{2}.$ (A.4)
The computation simplifies by employing that the noise is invariant under the rotations.
To simplify notation we remove the tilde and abbreviate the coefficients in the obvious way. Then $h_{\alpha}$ satisfies the evolution equations
$\partial_{t}h_{1}=2\Gamma_{3}(\partial_{x}h_{1})(\partial_{x}h_{2})+\tfrac{1}{2}D_{1}\partial_{x}^{2}h_{1}+\sigma_{1}\xi_{1},$
$\partial_{t}h_{2}=\Gamma_{1}(\partial_{x}h_{1})^{2}+\Gamma_{2}(\partial_{x}h_{2})^{2}+\tfrac{1}{2}D_{2}\partial_{x}^{2}h_{2}+\sigma_{2}\xi_{2}.$ (A.5)
There are still seven parameters and the goal now is to reduce their number. For this purpose we rescale the fields by choosing units of time, space, and amplitudes, through setting
$h_{\alpha}(x,t)=a_{\alpha}\tilde{h}_{\alpha}(x/\ell,t/\tau)\quad\mathrm{or}\quad\tilde{h}_{\alpha}(x,t)=(a_{\alpha})^{-1}h_{\alpha}(\ell x,\tau t)$ (A.6)
with adjustable coefficients $a_{\alpha},\ell,\tau$. Inserting in (A), one obtains
$\partial_{t}\tilde{h}_{1}=2\frac{a_{2}\tau}{\ell^{2}}\Gamma_{3}(\partial_{x}\tilde{h}_{1})(\partial_{x}\tilde{h}_{2})+\frac{\tau}{2\ell^{2}}D_{1}\partial_{x}^{2}\tilde{h}_{1}+\frac{\sqrt{\tau}}{a_{1}\sqrt{\ell}}\sigma_{1}\xi_{1},$
$\partial_{t}\tilde{h}_{2}=\frac{a_{1}^{2}\tau}{a_{2}\ell^{2}}\Gamma_{1}(\partial_{x}\tilde{h}_{1})^{2}+\frac{a_{2}\tau}{\ell^{2}}\Gamma_{2}(\partial_{x}\tilde{h}_{2})^{2}+\frac{\tau}{2\ell^{2}}D_{2}\partial_{x}^{2}\tilde{h}_{2}+\frac{\sqrt{\tau}}{a_{2}\sqrt{\ell}}\sigma_{2}\xi_{2}.$ (A.7)
The parameters $a_{1},\ell,\tau$ are chosen such that
$\frac{\tau}{\ell^{2}}D_{1}=\frac{\tau\sigma_{1}^{2}}{a_{1}^{2}\ell},\quad\frac{\tau}{\ell^{2}}D_{2}=1,\quad\frac{\sqrt{\tau}}{a_{2}\sqrt{\ell}}\sigma_{2}=1.$ (A.8)
Then the prefactor of $\tfrac{1}{2}\partial_{x}^{2}\tilde{h}_{1}$ turns into $(D_{1}/D_{2})$. Inserting this choice in the nonlinear term yields the successive coefficients of the nonlinearities,
[…]
Setting $a_{2}\Gamma_{2}=D_{2}$, the factor in front of the square bracket equals 1. Finally, setting
$(D_{1}/D_{2})=T,~{}(\Gamma_{3}/\Gamma_{2})=X,~{}(\sigma_{1}^{2}D_{2}\Gamma_{1}/\sigma_{2}^{2}D_{1}\Gamma_{2})=Y$ (A.10)
results in Eq. (1.4), as claimed.
In Sections 2.7 and 3.1 the following identity is used. We start from
$\partial_{t}h_{1}$ $=2\mathsfit{b}X(\partial_{x}h_{1})(\partial_{x}h_{2})+\tfrac{1}{2}\mathsfit{d}T\partial_{x}^{2}h_{1}+\sqrt{\mathsfit{d}T}\xi_{1},$ (A.11)
$\partial_{t}h_{2}$ $=\mathsfit{b}Y(\partial_{x}h_{1})^{2}+\mathsfit{b}(\partial_{x}h_{2})^{2}+\tfrac{1}{2}\mathsfit{d}\partial_{x}^{2}h_{2}+\sqrt{\mathsfit{d}}\xi_{2}.$
Setting $a_{1}=\mathsfit{d}/\mathsfit{b}$, $a_{2}=\mathsfit{d}/\mathsfit{b}$, $\ell=\mathsfit{d}^{2}/\mathsfit{b}^{2}$, $\tau=\mathsfit{d}^{3}/\mathsfit{b}^{4}$, the transformed fields satisfy
th1=2X(xh1)(xh2)+12Tx2h1+Tξ1,th2=Y(xh1)2+(xh2)2+12x2h2+ξ2.
[…]
Acknowledgments. HS thanks Jacopo De Nardis and Harvey Weinberger for most instructive discussions. AD, MK and HS thank the VAJRA faculty scheme (No. VJR/2019/000079) from the Science and Engineering Research Board (SERB), Department of Science and Technology, Government of India. AD and MK acknowledges the Department of Atomic Energy, Government of India, for their support under Project No. RTI4001.